OpenStudy (anonymous):

Use the fundamental trigonometric identities to simplify the expressions. Show all steps for full credit. HELP ME PLEASE 1.sec^2 x-tan^2 x= 2.cos^2 x(1+tan^2 x)= 3.secxcos(-x)-sin^2 x= 4.sec^2 x-1 / sin^2 x=

OpenStudy (anonymous):

first one is 1

OpenStudy (anonymous):

second one is 1

OpenStudy (anonymous):

the first one : 1/cos^2 -sin^2x/cos^2x= (1-sin^2x)/cos^2x= cos^2x/cos^2x=1

OpenStudy (anonymous):

second one: cos^2x((cos^2x + sin^2x)/cos^2x)=cos^2x *1/cos^2x=1

OpenStudy (anonymous):

thanks. :)

OpenStudy (anonymous):

the third one : 1/cosx * cos(x) -sin^2x = 1-sin^2x= cos^2x

OpenStudy (anonymous):

the last one: sec^2 x-1 / sin^2 x= (1/cos^2x -1)/sin^2x= ((1-cos^2x)/cos^2x)/sin^2x=(sin^2x/cos^2x)/sin^2x= 1/cos^2x= sec^2x

OpenStudy (jdoe0001):

\(\bf sec(x)cos(-x)-sin^2(x)\\ \quad \\ \textit{check you identities}\quad cos(-x)=cos(x)\qquad sec(x)=\cfrac{1}{cos(x)}\\ \quad \\ sec(x)cos(-x)-sin^2(x)\implies \cfrac{1}{\cancel{cos(x)}}\cdot \cancel{cos(x)}-sin^2(x)\\ \quad \\ 1-sin^2(x)\implies \square ?\)

OpenStudy (anonymous):

hope you understood sweetie!

OpenStudy (anonymous):

i do . thanks :)